A vacuum gage connected to a tank reads 30 kPa at a location where the barometric reading is 755 mm Hg. Determine the absolute pressure in kPa in the tank


1
Expert's answer
2021-08-12T11:02:25-0400

Vacuum gauge pressure= 30 KPa

Barometer reading =755 mm Hg

We know that the barometer always reads atmospheric pressure a given situation.

So atmospheric pressure is equal to 755 mm Hg.

We know that "P= ρ \times g \times h"

Density of Hg = 13600 kg/m3

"P=13600 \times 9.81 \times 0.755=100.72\; kPa"

Absolute pressure=atmospheric pressure + gauge pressure

But here given that 30 kPa is a Vacuum pressure, so we will take it as negative.

Absolute pressure=atmospheric pressure + gauge pressure

Absolute pressure "=100.72 - 30 =70.72\; kPa"

Answer: 70.72 kPs